OpenStudy (anonymous):

-6 less than or equal to 2x+6 less then or equal to 6

OpenStudy (anonymous):

welcome to open study please read the code of conduct http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

\[-6\leq 2x+6\leq 6\] subtract \(6\) all the way across then divide all by \(2\)

OpenStudy (anonymous):

-6 \[-6 \le 2x + 6 \le 6 \] \[-6-6\le2x + 6 - 6 \le 6-6\] \[-12\le2x \le0\] \[-6\le x \le 0 \]

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

yes, thank you